What happens to the rate of reaction involving CO, if the concentration is doubled and the reaction is:



first-order in CO? The new rate is equal to the old rate multiplied by what constant?





half-order in CO? The new rate is equal to the old rate multiplied by what constant?



the tolerance is +/-2%





inverse first-order in CO? The new rate is equal to the old rate multiplied by what constant?
1
Expert's answer
2016-11-03T08:12:16-0400
new rate = 2*old rate
new rate = √2*old rate
new rate=0.5*old rate